Find the equation (formula) of a sphere with radius r and center C(h, k, l) and show that x2 + y2 + z2 - 6x + 2y + 8z - 4 = 0 is an equation of a sphere. Also, find its center and radius


1
Expert's answer
2022-04-27T09:18:13-0400

"x^2 + y^2 + z^2 - 6x + 2y + 8z - 4 = 0 \\
x^2-6x+9-9+y^2+2y+1-1+z^2+8z+16-16-4=0 \\
x^2-6x+9 = (x-3)^2\\
y^2+2y+1 = (y+1)^2\\
z^2+8z+16 = (z+4)^2 ,then\\
(x-3)^2 - 9 +(y+1)^2 - 1+(z+4)^2-16-4 = 0\\
(x-3)^2 +(y+1)^2 +(z+4)^2 = 30\\
(x-a)^2 +(y-b)^2 +(z-c)^2 = R^2, then\\
a = 3,b=-1,c=-4, R^2 = 30"

Answer:

Equation of a sphere: "(x-3)^2 +(y+1)^2 +(z+4)^2 = 30\\"

Sphere center С(3, -1, -4)

Radius "\sqrt{30}"